For the following three vectors, what is 2 C.(3 A
Answer #1

2\,\vec{C}\,.\,(3\vec{A}\times\vec{B})=6\vec{C}\,.\,(\vec{A}\times\vec{B})

(\vec{A}\times\vec{B})=\begin{vmatrix} \hat{i} & \hat{j} &\hat{k} \\ 3.00&2.00 &-3.00 \\ -3.00&4.00 &3.00 \end{vmatrix}

(\vec{A}\times\vec{B})=\hat{i}(6+12)-\,\hat{j}(9-9)+\hat{k}(12+6)

(\vec{A}\times\vec{B})=18\hat{i}+18\hat{k}

6\vec{C}.(\vec{A}\times\vec{B})=6(7\hat{i}-8\hat{j})\,.\,(18\hat{i}+18\hat{k})

6\vec{C}.(\vec{A}\times\vec{B})=6\times7\times18 =756 (since \hat{i}.\hat{i}=1 ,\hat{i}.\hat{k}=0 ,\hat{j}.\hat{i}=0 , \hat{j}.\hat{k}=0 )